# An engine not a camera: Measuring performative power of online search

### v1

@@ Acknowledgements
We are particularly grateful to everyone who installed our extension and participated in the study; without you such a project would not have been feasible. Further, the authors would like to thank Jonathan Williams for the design of the project website, Alejandro Posada for producing the onboarding video and helping with the logo, Telintor Ntounis for assistance in setting up the server infrastructure, and Ana-Andreea Stoica, André Cruz and Jiduan Wu for feedback on the manuscript.
Appendix A Causal model
To support the discussion of our causal claim, consider the simplified diagram of how a user navigates to a website illustrated in Figure 7. The random variable $U$ denotes a user request and $T$ indicates which website is being visited in response. The user either navigates to the website via Google search (gray box), or they navigate to the website on some alternative way. This can be by entering the url directly, or using a different search service. Naturally, the arrangement $A$ only impacts the outcome $T$ if the user uses Google search, otherwise $A$ does not have any influence on the outcome. For a search query, the user query leads to an arrangement of content shown on the website, and the arrangement mediates the click. The setup in Section 3 focuses on performative power within the grey box, and Section 6 extends the claims to the full graph.
@@ Appendix A Causal model
Figure 7: Simplified causal graph illustrating the use of an online service. A web request leads to the visit of a website, potentially mediated by Google search.
Appendix B Proof of Theorem 1
@@
Using the definition of the performativity gap and the definition of the $L_{1}$ norm we have that
$\mathrm{PP}$ $=\sup_{a\in\mathcal{A}}\;\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\operatorname*{\mathbb{E}}|\mathrm{1}[C_{q}(a_{0})=c_{i}]-\mathrm{1}[C_{q}(a)=c_{i}]|$
$\geq\sup_{a\in\mathcal{A}}\;\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\left|\mathrm{Pr}[C_{q}(a_{0})=c_{i}]-\mathrm{Pr}[C_{q}(a)=c_{i}]\right|$
$\geq\sup_{a\in\mathcal{A}}\;\Big{|}\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathrm{Pr}[C_{q}(a_{0})=c_{i}]-\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathrm{Pr}[C_{q}(a)=c_{i}]\Big{|}$
$=\sup_{a\in\mathcal{A}}\delta^{i}(a)$
where the performativity gap is defined with respect to the set of queries $\mathcal{Q}$.
Appendix C Additional details on the study
C.1 Aggregate click statistics
In Figure 8 we show aggregate statistics over the clicks collected. In Figure 9 we provide aggregate statistics over the user base, collected through the onboarding form.
@@ C.1 Aggregate click statistics
Figure 9: Aggregate user statistics collected from the 85 participants with the onboarding form. Age distribution (left) and language in which they consume online search (right). This is the data we have about the demographics of our participants.
C.2 On the external validity of the Powermeter design
In our work we have demonstrated the large effect of arrangements on clicks for the participants of our study. We also discussed the robustness of our quantitative findings to variations in the sample of queries using bootstrapped confidence intervals.
Now discuss how the quantitative findings from the Powermeter experiment can be extrapolated to claims about a broader population. The key step is to argue that the observed sample of search queries is sufficiently representative for the population of interest. To this end, we note that the Powermeter design has the intrinsic property that the observed queries of an individual users follows their natural interaction with the platform. In turn, the representativeness of the user base has to be considered on a case-by-case basis, depending on the population under investigation. In technical terms, our sample of queries corresponds to a single-stage cluster sample of $\mathcal{Q}$. This is a commonly encountered sampling methodology in human subject studies.
Unbiasedness.
Our cluster sample estimates are unbiased, if the sample of users is also unbiased. Propensity score reweighting Stuart et al. (2011) can help to adjust estimates. To support such adjustments, Powermeter collects minimal demographic statistics about the users (see Figure 9). Further, this data is linked with collected clicks so the data can be reweighted appropriately.
Statistical error bounds.
For assessing the statistical error bounds of empirical estimates obtained with cluster sampling, it has to be taken into account to what extent the sample $S$ represent the variations in the population $\mathcal{Q}$. See (e.g., Rutterford et al., 2015; Abadie et al., 2022) for a reference. At the high level, assume there are $k$ types of users that respond differently to an arrangement. If all these user types are proportionally represented in the sample $S$, then the statistical uncertainty of our measurements scales with the number of queries as
$\delta^{1}(a)\geq\delta_{S}^{1}(a)-\mathcal{O}(1/\sqrt{|Q_{S}|}).$ (5)
In turn, if there is potential variation in the sampling of these types this also need to be taken into account. As long as variations across search queries dominate variations across users, the number of queries dominates the errors of the Powermeter estimates, but a correction term in (5) is necessary. Formally establishing these trade-offs, is an interesting step for future work.
Related to our findings, we conjecture that in the context of online search intra-cluster variation is more pronounced than inter-cluster variation. Meaning that the effect of position varies more across different type of search queries (whether it is an exploratory search for a suitable restaurant in a remote city, or a search for a specific research paper) than it does across users. This assessment might change is the consideration is restricted to specific types of queries.
C.3 Project website
The project website (https://powermeter.is.tue.mpg.de/) provides the entry point to the experiment. On the website users were provided with details about the experiment, instructions for how to participate, how to install the extension and information about how their data is used. We provide a list of frequently asked questions, as well as a instruction video. See Figure 10 for a screenshot. When clicking ‘participate’ the user’ is navigated to the extension in the Chrome store. We did not list the extension publicly to make sure the users have visited the project website before participating.
C.4 Onboarding form
@@
Upon installation of the extension the user is automatically navigated to the onboarding form, as illustrated in Figure 11. Filling the form is voluntary and answers are binned to only provide coarse grained information and no personally identifiable information. The information serves to gather some statistics about the demographics of the participants, but also to potentially debug different languages and website versions. Typically coarse-grained information about the statistics is sufficient to argue about qualitative properties of the population, e.g., for the sake of propensity adjustment.
C.5 Code
The extension is implemented in Java Script using the Google Manifest V3 framework. The extension is available through the Google Chrome store. The current version of the extension is v1.3. The experiment in this version is deactivated through a flag in the code. It is no longer maintained to keep up with recent updates on Google search. However, it is still available and the source code can be inspected using the Chrome DevTools.
Figure 10: Project website. Screencapture form May 29, 2024.
Figure 11: Onboarding form. Screencapture form May 29, 2024.
C.6 Extension policy
Below we provide the content of our extension policy. It is available through the Chrome store and published on our website:
[…] are committed to protect your data and to respect your privacy. This privacy statement explains the reason for the collection and processing of your data in the context of the Powermeter Research Experiment. It explains what data is collected, the way in which it is collected, handled, and protected, how that information is used.
Why and how do we process your data?
[…] collects and further processes your data to conduct the Powermeter Research Experiment (before, during and after) and to process your registration in that experiment. Powermeter is a Chrome web-browser extension (i.e. a software application adding functionality to a web browser) designed exclusively for research purposes and distributed through the Google Chrome Store. By installing the extension, users agree to participate in a research experiment. Powermeter performs random modifications to the display or positioning of search results in popular search engines to investigate the causal effect of display and positioning on user click behavior. Powermeter does not modify individual search results or links. It only rearranges ranking, positioning, and display of search results. It does not change, modify, distribute, or otherwise alter any part of the search engines’ software it operates on.
Powermeter does not collect personal data. It indexes users within the population through installation time-stamps and then counts the number of clicks occurring on search engines associated with different arrangements of content as inputted by each user. The indexing of users is necessary to understand if clicks are coming from different or from the same user and for the evaluation of statistics across different demographics. Specifically, Powermeter collects the following non-personal data: (i) the type of search engine used (Google or Bing); (iii) the version of the extension installed; (iii) the index of the search page where the click occurs; (iv) the user index; (v) the number and relative positioning of ads and specialised search results present on the page the click occurs; (vi) the relative positioning of the clicked item vis-à-vis related items object of interest; (vii) the timestamp of when the click occurred.
Beyond that, Powermeter never stores users’ search query, users’ identifiers (such as names, emails, IP addresses, etc.), nor records any information regarding the content of the links clicked over by users or the url of the landing page or any other data related to users’ browsing activity. The use of information received from Google APIs will adhere to the Chrome Web Store User Data Policy, including the Limited Use requirements.
At the time of installation, Powermeter will present users the option to answer a few multiple-choice questions concerning their gender, country, age, and language. Responding to these questions is facultative and collected answers will only be used to evaluate and compare aggregate click statistics across demographics.
Collected data is securely transmitted (using https protocol) and stored in a database on a secure and protected webserver [..]. Communication between the research team and adherents to the experiment will only occur through in-browser one-way (from the extension to the extension user) pop-up messages. Members of the research team will not be able to individually contact adherent users as they neither know who they are, nor are aware of their contact details.
The processing operations are linked to the organisation, functioning, and management, and follow-up to the Powermeter Research Experiment. Processing operations are limited to what is strictly necessary for the management and functioning of the Powermeter Research Experiment. […] Consequently, by downloading and installing the extension in the Chrome store you are giving us your explicit consent to:
  • Collect, store, and process your numeric installation time-stamp index;
  • Collect, store, and process your clicks associated with different arrangements of content on Google.com, Google.de, Google.ch and Bing.com;
  • Collect, store, and process the answer(s) you will voluntarily disclose concerning your gender, age, location, and language in the onboarding form;
  • Receive pop-up messages for any relevant communication related to the management, functioning, and termination of the Powermeter Research Experiment and to receive the messages hard coded within the extension. Data collection does only begin after Powermeter has been successfully installed.
Your consent can be withdrawn for the future at any time by uninstalling the extension. Your data will be used for the Powermeter Research Experiment only. Your data will not be shared, distributed, or repurposed.
Which personal data do we collect and further process?
No personal data will be collected or processed.
How long do we keep your data?
Your non-personal data will be only kept for the time strictly necessary to fulfil the purpose of collection or further processing.
All data related to the organisation and management of the Powermeter Research Experiment (this includes the information given during the registration, before, and during the experiment) will be deleted after the termination of the research project and within six months at the latest after the last action relating to the research project occurred.
How do we protect and safeguard your data?
All data are collected in electronic format, encrypted as described in Section n. 2, and stored on a secure and protected webserver […]. Access to the server is restricted and controlled.
Who has access to your data and to whom is it disclosed?
Access to your data is provided […] staff responsible for carrying out this processing operation and to other vetted and authorised researchers according to the “need to know” principle. The data we collect will neither be shared, nor transferred to any third party, and will not be repurposed.
What are your rights and how can you exercise them?
As an adherent user you are given the prerogative to interrupt the collection and processing of your data at any time by uninstalling the extension. The withdrawal of your consent will not affect the lawfulness of the processing carried out before you have withdrawn the consent.

### v2

@@ Acknowledgements
We are particularly grateful to everyone who installed the extension and participated in our study; without you such a project would not have been feasible. Further, the authors would like to thank Jonathan Williams for the design of the project website, Alejandro Posada for producing the onboarding video and helping with the logo, Telintor Ntounis for assistance in setting up the server infrastructure, and Ana-Andreea Stoica, André Cruz and Jiduan Wu for feedback on the manuscript. We would also like to thank two anonymous reviewers at NeurIPS who provided valuable feedback related to the presentation and framing of our work. Celestine Mendler-Dünner acknowledges the financial support of the Hector Foundation.
Appendix A Limitations and Broader Impact
We develop a flexible methodology, to provide insights into the power of digital platforms. We hope our framework can support future digital market investigations, complement and address some of the limitations of current antitrust tools. This could help make firms accountable for steering behavior on digital platforms of various kinds, and potentially support cases of anti-trust, consumer protection and abuse of dominance in digital markets. Compared to traditional tools, our approach offers a more straights forward way to integrate experimental insights with regulatory questions and requires fewer assumptions on the market dynamics. Furthermore, by providing a quantitative measurement, the methodology also allows to compare platforms and assess the effectiveness of potential remedies. That being said, instantiating the definition in a meaningful way is still at the discretion of the investigator and requires substantial domain knowledge. Further, fitting the approach within existing legal frameworks is an open question, we hope to further make this concrete in future work.
On a technical side, our design ensures that for any given users, the observed clicks follow a natural distribution. However, our participants form a convenience sample of online search users. Depending on the target of the investigation this might not be sufficient to argue for external validity of the quantitative insights. We provide statistics about our user base to support such a judgement in Figure 9. Further, we link these statistics with the observed clicks which can potentially help to adjust estimates using propensity score reweighting Stuart et al. (2011). However, more rigorously arguing about the external validity of our results in specific contexts is left for future work. But we expect the qualitative take-aways of our work to generalize beyond our study, and hope they can inform future modeling choices and problem statements around performativity in online search.
Lastly, we want to reiterate that our results for Bing should be taken with caution. While we designed our experiment to take most out of the available data, it is still a small sample of approximately $600$ search queries. Nevertheless, we did not want to withhold the results from the reader.
Appendix B Causal model
To support the arguments about composability of performative power in the discussion, consider the simplified causal diagram of how a user navigates to a website illustrated in Figure 7. The random variable $U$ denotes a user request and $T$ indicates which website is being visited in response. The user either navigates to the website via Google search (gray box), or they navigate to the website on some alternative way. This can be by entering the url directly, or using a different search service. Naturally, the arrangement $A$ only impacts the outcome $T$ if the user uses Google search, otherwise $A$ does not have any influence on the outcome. For every performed search query on Google, the user is exposed to an arrangement of content, and the arrangement mediates the click. Under appropriate independence assumptions the overall effect decomposes into two factors: the effect on outgoing traffic from Google search (gray box) and the overall fraction of relevant clicks being mediated by Google search. A formal treatment of this decomposition property under more general conditions is an interesting question for future work.
@@ Appendix B Causal model
Figure 7: Causal graph of online search users. A web request leads to the visit of a website, partially mediated by Google search.
Appendix C Proof of Theorem 1
@@
Using the definition of the performativity gap and the definition of the $L_{1}$ norm the proof is a direct consequence of
$\mathrm{PP}$ $=\sup_{a\in\mathcal{A}}\;\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\operatorname*{\mathbb{E}}|\mathrm{1}[C_{q}(a_{0})=c_{i}]-\mathrm{1}[C_{q}(a)=c_{i}]|$ (4)
$\geq\sup_{a\in\mathcal{A}}\;\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\left|\mathrm{Pr}[C_{q}(a_{0})=c_{i}]-\mathrm{Pr}[C_{q}(a)=c_{i}]\right|$ (5)
$\geq\sup_{a\in\mathcal{A}}\;\Big{|}\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathrm{Pr}[C_{q}(a_{0})=c_{i}]-\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathrm{Pr}[C_{q}(a)=c_{i}]\Big{|}$ (6)
$=\sup_{a\in\mathcal{A}}\delta^{i}(a)$ (7)
where the performativity gap is defined with respect to the set of queries $\mathcal{Q}$. In words, Theorem 1 formalizes the idea that the average effect of an arrangement change on an individual query $q\in\mathcal{Q}$ can be bounded by the average aggregate statistics across queries.
Appendix D Additional details on the study
D.1 Aggregate click statistics
In Figure 8 we show aggregate statistics over the clicks collected. In Figure 9 we provide aggregate statistics over the user base. The latter information was collected through the onboarding form.
@@ D.1 Aggregate click statistics
Figure 9: Aggregate user statistics collected from the 85 participants with the onboarding form. Age distirbution (left) and language in which they consume online search (right). That’s all the data we have about the demographics of our participants.
D.2 Project website
The project website provides the entry point to the experiment. Users where provided with details about the experiment, instructions for how to participate, and information about data usage. See Figure 10 for a screenshot.
Figure 10: Project website. URL and institution names are removed for the sake of anonymity.
D.3 Onboarding form
@@
Upon installation of the extension the user is navigated to the onboarding form, as illustrated in Figure 11. Providing the information is not mandatory and answers are binned to only provide coarse grained information and no personally identifiable information. The main purpose of the information serves debugging different languages and website versions.
Figure 11: Onboarding form.